find the value for k so that the remainder for (x^3 + 2x^2 - kx + 5)divided by (x - 2) is 3.
Answer by stanbon(75887)

find the value for k so that the remainder for (x^3 + 2x^2 - kx + 5)divided by (x - 2) is 3.
-------------------
Use synthetic division; the remainder must be f(2)=3
----------------------------
2)....1....2....-k....5
.......1....2...-k+4..|.13 -2k+8+5
=================
EQUATION:
-2k+13 = 3
-2k = -10
k = 5
==============
Cheers,
Stan H.